a bomb explodes in air when it has a horizontal a speed of 100 kilometre per hour it break into two piece a b of mass ratio 1:2 if a goes vertically up at a speed of 400 kilometre per hour find the velocity of be immediately after die explosion

Grade:11

2 Answers

Vikas TU
7 years ago
From the momentum conservation,
(3m*100)icap = (m*400)jcap + 2m*v
Further solving we get,
2v = 300icap – 400jcap
v = 150i – 200j Km/hr.
Arun
4 years ago
From the momentum conservation,
(3m*100)icap = (m*450)jcap + 2m*v
Further solving we get,
2v = 300icap – 450jcap
v = 150i – 200j Km/hr.